OpenStudy (anonymous):

Find the derivative of : y=log(base 3)x^2 show work thanks

OpenStudy (asnaseer):

Using log rules:\[y=log_3x^{2}=\frac{\ln x^{2}}{\ln 3}=\frac{2\ln x}{\ln 3}\]\[dy/dx=\frac{2}{x\ln 3}\]

OpenStudy (anonymous):

I found the derivative of the numerator and got (2/x)/(ln3)...... how does that turn into the final answer you got?

OpenStudy (asnaseer):

\[(2/x)/\ln 3=\frac{2}{x} * \frac{1}{\ln 3}=\frac{2}{x\ln 3}\]

OpenStudy (anonymous):

thanks